A weak acid has a pKa of 7.89. calculate the pH of a solution in which the ratio of the concentration of base to acid (base/acid) equals 2.92
1
Expert's answer
2020-05-15T10:05:39-0400

pH=pKa+"\log{\frac{base}{acid}}"


pH=7.89+"\log2.92" =7.89+0.47=8.36